# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

### v1

@@ 4 Conformal methods for box coordinates
A key modelling decision in CP is the choice of scoring function, $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, for computing nonconformity scores (see Algorithm 1). We consider three choices of scoring function and subsequent PI construction for every box coordinate $k\in\{1,\dots,m\}$. Additional implementation details can be found in Appendix 0.B.
@@
Standard conformal (Box-Std). We firstly consider the standard approach of using regression residuals as scores [52], i.e., $s(\hat{f}_{B}(X),Y)=|\hat{c}^{k}-c^{k}|$. The resulting conformal PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{q}^{k}_{B}]$, where $\hat{q}^{k}_{B}$ denotes the computed conformal quantile for the $k$-th coordinate. While simple, this construction only allows for non-adaptive, fixed-width PIs. Andéol et al. [14] use this approach to construct their one-sided intervals.
@@
Conformal ensemble (Box-Ens). In order to produce more adaptive PIs, we next consider using normalized residual scores [30] of the form $s(\hat{f}_{B}(X),Y)=|\hat{c}^{k}-c^{k}|/\hat{\sigma}(X)$, where $\hat{\sigma}$ is some choice for a heuristic uncertainty estimate (i.e., without guarantees) produced by the underlying object detector. The resulting conformal PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B}]$. By incorporating model uncertainty, the constructed PIs can be re-scaled individually per coordinate, adapting their magnitudes at test time. We may interpret this as a way of empirically conditioning on the given test sample. We employ an ensemble of object detectors and quantify $\hat{\sigma}$ as the standard deviation of the ensemble’s box coordinate predictions [27]. A joint coordinate prediction $\hat{c}^{k}$ is obtained via confidence-weighted box fusion [54].
@@
Conformal quantile regression (Box-CQR). As a second adaptive method, we fit the score and interval construction of Conformal Quantile Regression (CQR) [48] to our setting. Additional regression heads $\hat{Q}_{B}$ are added to the object detector alongside $\hat{f}_{B}$ and trained with a quantile loss to produce lower and upper quantile predictions $\hat{Q}_{\alpha_{B}/2}$ and $\hat{Q}_{1-\alpha_{B}/2}$ for the bounding box coordinates. [^1] Note that these predictions are not the same as conformal quantiles. Under regularity conditions, these will converge asymptotically to the true conditional quantiles [23, 11] and achieve target coverage $(1-\alpha_{B})$. Following CQR, we define the scores as $s(\hat{Q}_{B}(X),Y)=\max\{\hat{Q}_{\alpha_{B}/2}(X)-c^{k},\,c^{k}-\hat{Q}_{1-\alpha_{B}/2}(X)\}$, and construct the conformal PIs as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{Q}_{\alpha_{B}/2}(X_{n+1})-\hat{q}^{k}_{B},\,\,\hat{Q}_{1-\alpha_{B}/2}(X_{n+1})+\hat{q}^{k}_{B}]$. The obtained PI ensures adaptivity through the use of quantile predictions, which will differ in their distance relative to the mean coordinate prediction $\hat{c}^{k}_{n+1}$ for each sample.
@@ 4.1 Multiple testing correction
CP can also be interpreted from a hypothesis testing perspective [62, 53]. A null hypothesis for the $k$-th test sample coordinate is formed as $H_{0}:c^{k}_{n+1}=c^{k}$ for some candidate value $c^{k}\in\mathbb{R}$. Leveraging the nonconformity scores of the calibration samples as an empirical null distribution, a p-value $p(c^{k})$ is computed as the fraction of samples which conform worse than $s(\hat{c}^{k}_{n+1},c^{k})$. The value $c^{k}$ is included in the prediction set — i.e., we cannot reject $H_{0}$ — if $p(c^{k})>\alpha_{B}$. Conformalizing a set of $m$ coordinates separately and thus running $m$ nonconformity tests in parallel gives rise to multiple testing issues. Without correction, the guaranteed coverage reduces to $(1-m\,\alpha_{B})$ (see subsection 0.A.3).
A naive correction can be achieved using Bonferroni [64], since choosing $\alpha_{B}^{\prime}=\alpha_{B}/m$ will satisfy target coverage. However, the Bonferroni correction is overly conservative under positive dependency of the individual hypothesis [63], which is reasonable to assume given that the coordinates parametrize an object’s bounding box jointly. In fact, Bates et al. [8] assert that a set of conformal p-values exhibits positive dependency structure a priori since they are jointly positive regression dependent on a subset [9].
We leverage an alternate procedure by Timans et al. [58], which exploits correlation structure among box coordinates for a less conservative correction. Their suggested algorithm $\max$-rank is an adaptation of the Westfall & Young [67] permutation correction suitable for the conformal setting (see subsection 0.A.4), permitting its integration into our approach. While the use of a simple $\max$-correction has been considered for CP [1, 10], $\max$-rank operates in the scale-invariant rank space and is thus robust. In addition, it requires less computational overhead than proposed copula-based corrections [57, 37].
@@ 5 Class label prediction sets
Given our multi-class setting, the object detector is at risk of incorrectly predicting an object’s class label. This raises issues for a direct application of class-conditional CP to the object’s bounding box at test time, since we require correctly selecting the conformal box coordinate quantiles $\hat{q}^{k}_{B},\,k\in\{1,\dots,m\}$ to construct PIs that satisfy Equation 3. This limits any provided nominal safety assurances to correctly classified objects only, i.e., we require $\hat{l}_{n+1}=l_{n+1}$ for any class $y\in\mathcal{Y}$.
In order to relax the above condition on the model’s classification abilities, we introduce an additional conformal step into our modelling pipeline which preceeds the conformal bounding box construction (see Figure 2). Specifically, we consider conformalizing the model’s classifier head $\hat{f}_{L}$ to produce class label prediction sets $\hat{C}_{L}(X_{n+1})$ which guarantee correct label containment. These are subsequently used to select our box coordinate quantiles, ensuring the validity of the provided guarantee in Equation 3 is broadened to also include any incorrectly classified objects.
We achieve this using a class-conditional CP approach on the class labels with a strict label coverage guarantee of 99% (i.e., setting $\alpha_{L}=0.01$), thus approximating the condition $\hat{l}_{n+1}=l_{n+1}$ by effectively ensuring $\hat{C}_{L}(X_{n+1})\ni l_{n+1}$. The resulting two-step sequential conformal approach maintains validity regardless of the object detector’s classification or bounding box regression performance – the incurred costs are reflected in the obtained prediction set and interval sizes only. Our experiments in subsection 6.3 highlight that even under full safety assurances, our approach provides practically actionable results. We follow with a description of the employed conformal label set method and considered baselines.
@@ 5.1 Conformal class thresholding (ClassThr)
We propose using a class-conditional variant of the prediction set classifier introduced by Sadinle et al. [50], based on a similar split conformal procedure as before (see Algorithm 2). Given our probabilistic classifier $\hat{f}_{L}$, we define the scoring function $s(\hat{f}_{L}(X),y)=1-\hat{\pi}_{y}(X)$ for every class $y\in\mathcal{Y}$ and obtain per-class conformal label quantiles $\hat{q}_{L}^{y}$. The conformal label prediction set for a new object to classify is then given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\}$ [^2] We invert the scores used by Sadinle et al. [50] to better reflect the notion of nonconformity, while being equivalent in procedure.. Importantly, class-conditional validity is ensured by comparing each class probability against its class-specific threshold on set inclusion.
The class-conditional guarantee stated in Equation 2 can now be provided equivalently for the object detector’s classifier head, i.e., we have that
@@
Impact on bounding box coverage guarantee. We enforce a class-conditional guarantee on the label prediction sets, since nominally requiring only a weaker marginal guarantee risks invalidating the subsequent class-conditional box coverage guarantee. If a class is systematically undercovered, we may fail to retrieve the correct box quantiles for some of its objects. Unless the class-specific label quantile is always dominated by some quantile associated with the labels present in the set, any undercoverage tendencies propagate down-stream.
The strength of the label guarantee impacts the box guarantee, since we perform both CP procedures in sequence. We observe that both guarantees are controlled in two distinct conformal procedures on the calibration data, and thus conditionally independent. It then follows for their sequential application that
$\begin{split}&\mathbb{P}(l_{n+1}\in\hat{C}_{L}(X_{n+1})\,\wedge\,\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))\,|\,l_{n+1}=y)\\ &=\mathbb{P}(l_{n+1}\in\hat{C}_{L}(X_{n+1})|l_{n+1}=y)\,\cdot\,\mathbb{P}(\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))|l_{n+1}=y)\\ &\hskip 150.00023pt\geq(1-\alpha_{L})(1-\alpha_{B})\quad\forall y\in\mathcal{Y}.\end{split}$ (5)
That is, a preceeding label coverage guarantee of $1-\alpha_{L}$ will nominally only assure subsequent box coverage of $(1-\alpha_{L})(1-\alpha_{B})$. In our experiments we are able to empirically approximate $(1-\alpha_{L})(1-\alpha_{B})\approx 1-\alpha_{B}$ by setting $\alpha_{L}=0.01$, thus alleviating the downstream coverage reduction. Equation 5 furthermore highlights that coverage trade-offs between objectives are possible according to sought-after assurances. For example, nominal box coverage of 90% can be achieved by both selecting $\alpha_{L}=0.05,\alpha_{B}=0.05$ and $\alpha_{L}=0,\alpha_{B}=0.1$.
@@ 5.2 Bounding box quantile selection
Obtained label prediction sets are utilized to select a conformal box coordinate quantile which provides a valid bounding box PI construction. A natural quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}\,\forall k\in\{1,\dots,m\}$, where $\hat{q}_{B}^{k,y}$ is the quantile of the $k$-th coordinate for class $y$. Using a $\max$-operator on the label set is a cautious but conservative approach, since each label in the set is regarded as equally likely. We thus incur an overcoverage cost in terms of wider PIs, which may perhaps be alleviated with a different strategy. However, we find even this simple approach to yield reasonably sized box intervals. A hypothesis testing motivation can be found in subsection 0.A.5.
@@
Top $*{}^{*}$ Single ${}^{\dagger}$ Small
Naive $*{}^{*}$ Small ${}^{\dagger}$ Small
@@
Table 1: Provided nominal coverage guarantees and expected empirical prediction set/interval sizes for the considered label prediction set methods on the basis of both correctly and incorrectly classified objects. $*{}^{*}$op and Naive provide a label guarantee ‘for free’ if $(1-\alpha_{L})$ is below the classifier’s accuracy level for each class. ${}^{\dagger}$op and Naive provide a box guarantee for correctly classified objects only. Naive may also satisfy both guarantees assuming practically unattainable perfect model calibration.
@@ 5.3 Label set baselines
We compare conformal class thresholding (ClassThr) to a series of alternative label prediction set constructions, whose nominal guarantees and empirically expected set sizes are outlined in Table 1.
Top singleton set (Top). We return label prediction sets that only consist of the highest probability class for every sample, i.e., $\hat{C}_{L}(X_{n+1})=\{y^{*}:\hat{\pi}_{y^{*}}(X_{n+1})=\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}(X_{n+1})\}$. The approach returns singleton sets and comes without nominal guarantees, and its empirical coverage relies fully on the classifier’s accuracy. The distinction to our initial condition $\hat{l}_{n+1}=l_{n+1}$, which we refer to as Oracle, is subtle: instead of ensuring correct quantile selection, we permit the use of potentially wrong quantiles to construct the box PIs.
Density level set (Naive). Assuming a perfectly calibrated classifier such that $\hat{\pi}_{y}(X)=\pi_{y}(X)\,\,\forall y\in\mathcal{Y}$, the optimal prediction set is provided by density level sets. That is, we collect all labels sorted by descending $\hat{\pi}_{y}$ until we reach probability mass $1-\alpha_{L}$. Under this assumption, prediction sets then also satisfy conditional coverage for any $X\in\mathcal{X}$ [50, 49]. While unattainable in practical settings where the classifier tends to be miscalibrated (i.e., $\hat{\pi}_{y}(X)\neq\pi_{y}(X)$), it can be considered a theoretically motivated extension of the Top baseline.
Full domain set (Full). We consider the label prediction set to be the full set of possible class labels for every sample, i.e., $\hat{C}_{L}(X_{n+1})=\mathcal{Y}$ and $|\hat{C}_{L}(X_{n+1})|=K$. In combination with our quantile selection strategy outlined below, this approach provides full label coverage for every sample, that is, it ensures $\alpha_{L}=0$. However, the cost of overly inflated label prediction sets is expected to propagate to subsequent bounding box PI construction.
We do not consider other popular conformal approaches for classification such as APS [49] or RAPS [5] since they aim at empirically improving conditional coverage under the requirements of a marginal guarantee – their advantages do not extend to a class-conditional setting as ours.

### v2

@@ 4 Conformal methods for box coordinates
A key modelling decision in CP is the choice of scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ to compute the required nonconformity scores (see § 0.A.1). We consider three choices of scoring function and PI construction for each box coordinate $k\in\{1,\dots,m\}$, which we outline next. Additional implementation details can be found in Appendix 0.B.
@@
Standard conformal (Box-Std). We firstly consider the standard approach of employing regression residuals $s(\hat{f}_{B}(X),\,Y)=|\hat{c}^{k}-c^{k}|$ as scores [54]. The resulting PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{q}^{k}_{B}]$, where $\hat{q}^{k}_{B}$ denotes the computed conformal quantile for the $k$-th coordinate. While straightforward, this construction only permits for non-adaptive, fixed-width intervals. Andéol et al. [15] use this approach to construct their one-sided intervals.
@@
Conformal ensemble (Box-Ens). In order to produce more adaptive intervals, we next consider normalized residual scores [31] of the form $s(\hat{f}_{B}(X),\,Y)=|\hat{c}^{k}-c^{k}|/\hat{\sigma}(X)$, where $\hat{\sigma}$ is some form of heuristic uncertainty estimate (i.e., without guarantees) obtained from the underlying model. The resulting conformal PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B}]$. By incorporating model uncertainty, the intervals can be re-scaled individually per coordinate to adapt their magnitude at test time. We can interpret this as an empirical conditioning on the particular test sample. We employ an ensemble of object detectors and quantify $\hat{\sigma}$ as the standard deviation of the ensemble’s box coordinate predictions [28]. A joint coordinate prediction $\hat{c}^{k}$ is obtained from the ensemble via confidence-weighted box fusion [56].
@@
Conformal quantile regression (Box-CQR). As an alternative adaptive method, we extend the approach of Conformal Quantile Regression (CQR) [50] to our setting. Additional regression heads $\hat{Q}_{B}$ are trained with a quantile loss alongside $\hat{f}_{B}$ to produce lower and upper quantile predictions $\hat{Q}_{\alpha_{B}/2}$ and $\hat{Q}_{1-\alpha_{B}/2}$ for the bounding box coordinates. Under regularity conditions, these predictors will asymptotically converge to the true conditional quantiles [24, 12], motivating their viability. Following CQR, we define the scores as $s(\hat{Q}_{B}(X),\,Y)=\max\{\hat{Q}_{\alpha_{B}/2}(X)-c^{k},\,c^{k}-\hat{Q}_{1-\alpha_{B}/2}(X)\}$, and construct the conformal PIs as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{Q}_{\alpha_{B}/2}(X_{n+1})-\hat{q}^{k}_{B},\,\,\hat{Q}_{1-\alpha_{B}/2}(X_{n+1})+\hat{q}^{k}_{B}]$. The obtained interval ensures adaptivity through the use of its quantile predictions, which will differ in their distance relative to the mean coordinate prediction $\hat{c}^{k}_{n+1}$ per sample.
@@ 4.1 Multiple testing correction
Applying CP to each of the box coordinates $k=1,\dots,m$ separately gives rise to multiple testing issues, since the conformal procedure can be interpreted from a hypothesis testing view as running $m$ permutation tests on nonconformity in parallel [65, 55]. This results in a guaranteed coverage of at most $(1-m\cdot\alpha_{B})$, as opposed to the desired rate of $(1-\alpha_{B})$ (see § 0.A.3 for further details).
The naive Bonferroni correction [67] offers a possible remedy, since choosing $\alpha_{B}^{\prime}=\alpha_{B}/m$ will satisfy target coverage. However, it is known to be overly conservative under positive dependency of the individual hypothesis [66], a reasonable assumption given that the coordinates parametrize an object’s bounding box jointly. In fact, Bates et al. [8] assert that a set of conformal p-values exhibits positive dependency structure a priori since they are jointly positive regression dependent on a subset [9]. We leverage an alternate procedure by Timans et al. [61], which exploits correlation structure among box coordinates for a less conservative correction. Their $\max$-rank procedure adapts the Westfall & Young [70] permutation correction to make it suitable for the setting of conformal prediction (see § 0.A.4). While the use of a $\max$-correction has been previously considered for CP [1, 11], $\max$-rank operates in the more robust scale-invariant rank space. In addition, it requires less compute than previously proposed copula-based testing corrections [60, 38].
@@ 5 Class label prediction sets
In practice, the object detector may incorrectly predict an object’s class label given our multi-class setting. This complicates a direct application of class-conditional CP to an object’s bounding box at test time, since we need to correctly select the conformal quantiles $\hat{q}^{k}_{B},\,k\in\{1,\dots,m\}$ to construct bounding box intervals that satisfy Eq. 3. This limits intially provided safety assurances to correctly classified objects only, i.e., those where $\hat{l}_{n+1}=l_{n+1}$ successfully match.
To alleviate this restrictive dependence on the model’s classification ability, our modelling pipeline in Fig. 2 introduces an additional conformal step which preceeds the bounding box construction. Specifically, we consider applying CP to the model’s classifier head $\hat{f}_{L}$ to first generate class label prediction sets $\hat{C}_{L}(X_{n+1})$ with a guarantee on label containment. These are subsequently used to select our box coordinate quantiles, ensuring the validity of our provided guarantees in Eq. 3 is broadened to even include incorrectly classified objects.
We achieve this using another class-conditional CP approach on the class labels with a strict label coverage guarantee of 99% (i.e., $\alpha_{L}=0.01$). Thus, we approximate the condition $\hat{l}_{n+1}=l_{n+1}$ by effectively ensuring $l_{n+1}\in\hat{C}_{L}(X_{n+1})$. The resulting two-step sequential approach maintains validity regardless of the object detector’s classification or bounding box regression performance. The incurred costs are reflected in the obtained prediction set and interval sizes only. Our experiments in § 6.3 demonstrate that even under these strong safety assurances, our approach provides actionably tight PIs. We follow with a description of the employed conformal label set method and related baselines.
@@ 5.1 Conformal class thresholding (ClassThr)
We propose using a class-conditional variant of the prediction set classifier introduced by Sadinle et al. [52], based on a similar conformal procedure as for the bounding boxes. Given our probabilistic classifier head $\hat{f}_{L}$, we define the scoring function $s(\hat{f}_{L}(X),\,y)=1-\hat{\pi}_{y}(X)$ for every class $y\in\mathcal{Y}$ to compute per-class conformal quantiles $\hat{q}_{L}^{y}$. The class label prediction set for a new object to classify is then given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\}$. The detailed procedure is given in Algorithm 2, deferred to § 0.A.2. Importantly, class-conditional validity is ensured by comparing each class probability against its class-specific threshold on set inclusion. The class-conditional guarantee stated in Eq. 2 can now be provided similarly for the object detector’s classification task as
@@
Impact on bounding box coverage guarantee. A class-conditional guarantee is enforced for the label prediction sets (Eq. 4), since only imposing the weaker marginal guarantee could invalidate the subsequent class-conditional box guarantee (Eq. 3). If for instance a class is systematically undercovered, we would fail to retrieve the correct box quantiles for some of its associated objects, propagating the undercoverage down-stream. Our approach instead enforces guarantees of equivalent strength. Observe that we perform two distinct conformal procedures in sequence, rendering the coverage guarantees conditionally independent. Thus the down-stream coverage effect for their two-step application is that
$\begin{split}&\mathbb{P}\left(l_{n+1}\in\hat{C}_{L}(X_{n+1})\,\wedge\,\bigcap_{k=1}^{m}\left(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1})\right)\,|\,l_{n+1}=y\right)\\ &=\mathbb{P}\left(l_{n+1}\in\hat{C}_{L}(X_{n+1})\,|\,l_{n+1}=y\right)\,\cdot\,\mathbb{P}\left(\bigcap_{k=1}^{m}\left(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1})\right)\,|\,l_{n+1}=y\right)\\ &\hskip 180.00027pt\geq(1-\alpha_{L})(1-\alpha_{B})\quad\forall y\,\in\mathcal{Y}.\end{split}$ (5)
That is, a preceeding label coverage guarantee of $(1-\alpha_{L})$ will nominally only assure subsequent box coverage of $(1-\alpha_{L})(1-\alpha_{B})$. In our experiments we approximate $(1-\alpha_{L})(1-\alpha_{B})\approx(1-\alpha_{B})$ by setting $\alpha_{L}=0.01$, thus alleviating the down-stream coverage reduction [^2] As we observe in § 6.3, enforcing $\alpha_{L}=0$ leads to empirically inefficient box intervals.. Eq. 5 highlights that coverage trade-offs between objectives are possible depending on application-specific requirements. For example, a nominal box coverage of 90% can be achieved by choosing either $\alpha_{L}=0.05,\alpha_{B}=0.05$ or $\alpha_{L}=0,\alpha_{B}=0.1$.
@@ 5.2 Bounding box quantile selection
The obtained label prediction sets are subsequently used to select a valid box coordinate quantile for any interval construction following § 4. A natural quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}\,\forall k\in\{1,\dots,m\}$, where $\hat{q}_{B}^{k,y}$ is the quantile of the $k$-th coordinate for class $y$. Using a $\max$-operator on the label set is a valid but conservative approach, since all labels in the set are regarded as equally likely for every sample. Obtained box intervals thus tend to overcover, which could perhaps be alleviated with a different strategy, resulting in narrower PIs. Yet, we find even this straightforward selection to yield reasonably tight results. A hypothesis testing motivation for its use can be found in § 0.A.5.
@@
Top Single Small
Naive Small Small
@@
Table 1: Provided nominal coverage guarantees and expected empirical prediction set/interval sizes for the considered label prediction set methods on the basis of both correctly and incorrectly classified objects. Top and Naive provide a label guarantee ‘for free’ if $(1-\alpha_{L})$ is below the classifier’s accuracy level for each class. Top and Naive provide a box guarantee for correctly classified objects only. Naive may also satisfy both guarantees under practically unattainable perfect model calibration.
@@ 5.3 Label set baselines
We compare obtained label predictions sets via conformal class thresholding (ClassThr) to several reasonable alternatives, whose nominal guarantees and expected empirical set sizes are outlined in Tab. 1, and which we detail next.
Top singleton set (Top). We return label prediction sets that only consist of the highest probability class for every sample, i.e., $\hat{C}_{L}(X_{n+1})=\{y^{*}:\hat{\pi}_{y^{*}}(X_{n+1})=\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}(X_{n+1})\}$. This approach returns singleton sets, is void of nominal guarantees, and its empirical coverage relies fully on the classifier’s accuracy. The distinction to our initial condition $\hat{l}_{n+1}=l_{n+1}$, which we refer to as Oracle, is subtle: instead of ensuring correct quantile selection, we permit the use of potentially wrong quantiles to construct the box intervals.
Density level set (Naive). Assuming a perfectly calibrated classifier such that $\hat{\pi}_{y}(X)=\pi_{y}(X)\,\,\forall y\in\mathcal{Y}$, the optimal prediction set is provided by density level sets. That is, we collect all labels sorted by descending $\hat{\pi}_{y}$ until we reach probability mass $(1-\alpha_{L})$. Under this assumption, prediction sets will also approach conditional coverage for any $X\in\mathcal{X}$ [52, 51]. While unattainable in practical settings where the classifier tends to be miscalibrated (i.e., $\hat{\pi}_{y}(X)\neq\pi_{y}(X)$), it can be considered a theoretically motivated extension of the Top baseline.
Full domain set (Full). We consider taking the full set of possible class labels per sample, thus $\hat{C}_{L}(X_{n+1})=\mathcal{Y}$ and $|\hat{C}_{L}(X_{n+1})|=K$. In combination with our quantile selection strategy this approach guarantees label coverage conditionally per sample, i.e., it ensures $\alpha_{L}=0$. However, this comes at the cost of overly inflated label sets whose size is expected to propagate to the box intervals.
We do not consider other popular conformal approaches for classification such as APS [51] or RAPS [5] since they aim to empirically improve conditional coverage under the requirements of a marginal guarantee – these advantages do not extend to a class-conditional setting as ours.